What is 5x + 2y =-22 9x-2y=-62 solved
in the addition method ?

Answer by Math_Boss(45)

5x+2y=-22
9x-2y=-62
Add:
14x+0=-84
14x=-84
x=-6
Substitute x=(-6)
5(-6)+2y=-22
-30+2y=-22
2y=-22+30
2y=8
y=4
Ans:(-6, 4)